# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v1

@@ Abstract
This paper studies constrained Markov decision processes (CMDPs) with constraints against stochastic thresholds, aiming at safety of reinforcement learning in unknown and uncertain environments. We leverage a Growing-Window estimator sampling from interactions with the uncertain and dynamic environment to estimate the thresholds, based on which we design Stochastic Pessimistic-Optimistic Thresholding (SPOT), a novel model-based primal-dual algorithm for multiple constraints against stochastic thresholds. SPOT enables reinforcement learning under both pessimistic and optimistic threshold settings. We prove that our algorithm achieves sublinear regret and constraint violation; i.e., a reward regret of $\tilde{\mathcal{O}}(\sqrt{T})$ while allowing an $\tilde{\mathcal{O}}(\sqrt{T})$ constraint violation over $T$ episodes. The theoretical guarantees show that our algorithm achieves performance comparable to that of an approach relying on fixed and clear thresholds. To the best of our knowledge, SPOT is the first reinforcement learning algorithm that realises theoretical guaranteed performance in an uncertain environment where even thresholds are unknown.
@@ 1 Introduction
Reinforcement learning has enabled major advances in AI, such as AlphaGo silver2016mastering , autonomous vehicle control kendall2019learning , and LLM alignment ouyang2022training , in which an agent is trained to maximise its expected cumulative reward from the environment Sutton1998 . Markov decision processes (MDPs) are widely adopted in reinforcement learning, where actions and rewards depend only on the current state rather than on the entire history bellman1957markovian . In many applications, ensuring safety is of profound importance – for instance, autonomous vehicles are expected to avoid collisions and adhere to traffic regulations wen2020safe . This motivates constrained Markov decision processes (CMDPs), introducing safety constraints against thresholds into optimisation of the cumulative reward function altman1999constrained .
@@
  • The first approach to safe reinforcement learning agents in such uncertain environment. SPOT leverages a Growing-Window estimator sampling during the agent’s interactions with the environment to estimate the thresholds. Based on this estimator, SPOT employs the primal-dual scheme to integrate the constraints into the learnt agent. SPOT is compatible to two variants that (1) employs pessimistic thresholds to strictly enforce safety, and (2) applies optimistic thresholds to encourage exploration and boost performance. By systematically navigating the latent ‘margin’ between safety and performance without relying on any prior threshold information, our approach offers a dynamic balance between exploitation and exploration.
@@ 2 Preliminaries
An MDP characterises the sequential decision-making process of an agent, positioned in a state $s$ from a finite state $\mathcal{S}$ (with $|\mathcal{S}|=S$), which could select an action $a$ from a finite action space $\mathcal{A}$ (with $|\mathcal{A}|=A$) and then transition to another state $s^{\prime}\in\mathcal{S}$ according to a transition function $p_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$. The course of each episode starts at an initial state $s_{1}\in\mathcal{S}$ and continues for $H\in\mathbb{N}$ steps, which is named as horizon. At each step $h\in[H]$, the agent receives a non-stationary reward $\tilde{r}_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ drawn from a distribution $\mathcal{R}$ with expectation reward $r_{h}(s,a)$, which means its underlying distribution may vary over time. The agent’s trajectory is determined by a policy $\pi$ consisting of decision rules $(\pi_{1},\ldots,\pi_{H})\in\Pi$, where each $\pi_{h}(\cdot|s)\in\Delta(\mathcal{A})$ is a probability distribution over the action space $\mathcal{A}$. Here, $\Delta(\mathcal{A})$ denotes the probability simplex over the action space $\mathcal{A}$. A Constrained MDP (CMDP) extends an MDP by incorporating $m$ constraints, each defined by a function $g_{i,h}:\mathcal{S}\times\mathcal{A}\to[0,1]^{m}$ associated with a threshold. These thresholds collectively form a vector $\alpha\in[0,H]^{m}$. In the non-stationary scenario, the immediate constraint $\tilde{g}_{i,h}(s,a)\in[0,1]$ is drawn from a distribution $\mathcal{G}_{i}$ with expectation $\mathbb{E}_{\mathcal{G}_{i}}[\tilde{g}_{i,h}(s,a)]=g_{i,h}(s,a)$ for every constraint $i\in[m]$. Overall, a CMDP can be represented by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},H,p,r,g,\alpha)$.
@@
capture the expected sum of $v$ from step $h$ onward. For brevity, let $V_{v}^{\pi}$ denote $V_{v,1}^{\pi}(s_{1})$, the expected return starting at the initial state. The primary objective is to find a policy that maximises the expected cumulative reward while satisfying all the constraints; formally,
@@
The reward regret measures the cumulative gap between the rewards obtained by the selected policies and the optimal policy $\pi^{\star}\in\Pi$, and the constraint violation measures how much the constraints are violated during the learning process. The objective of the agent is to select policies $\pi_{t}\in\Pi$ for each episode $t\in[T]$ such that both $\mathcal{R}_{T}(r)$ and $\mathcal{R}_{T}(g)$ grow sublinearly in the number of episodes $T$.
@@ 3.1 Unknown and Stochastic Environment in Terms of Thresholds
Motivated by the examples above, we propose training a CMDP in which thresholds are unknown, and must therefore be inferred through real-time sampling. Specially, for constraint $i\in[m]$, the threshold $\tilde{\alpha}_{i,h}(s,a)\in[0,H]$ for any state-action pair $(s,a)$ is drawn from a distribution $\mathcal{L}_{i}$ with support of $[0,H]^{\mathcal{S}\times\mathcal{A}}$, whose expectation is $\mathbb{E}_{\mathcal{L}_{i}}[\tilde{\alpha}_{i,h}(s,a)]=\alpha_{i,h}(s,a)$.
Algorithm 1 depicts how an agent interacts with the environment, when rewards, constraints and thresholds are all stochastic. At each episode $t\in[T]$, the agent chooses a policy $\pi_{t}$, while independently sampling reward $r^{t}\sim\mathcal{R}$, constraint $g_{i}^{t}\sim\mathcal{G}_{i}$ and threshold $\alpha_{i}^{t}\sim\mathcal{L}_{i}$ for every constraint $i\in[m]$. It is important to note that the agent has no prior knowledge of the reward, constraint, and threshold distributions, nor the transition function $p$.
@@
Also motivated by the examples, we note that the environment is not static – rewards, constraints, and thresholds change over time. To take these into account, we define that constraints, and thresholds are drawn from dynamical distributions $\mathcal{R}_{t}$, $\mathcal{G}_{t}$, $\mathcal{L}_{t}$, over $T$ episodes, instead of fixed distributions. Consequently, their expectations $\tilde{r}_{h}^{t}(s,a)$, $\tilde{g}_{i,h}^{t}(s,a)$, $\tilde{\alpha}_{i,h}^{t}(s,a)$ also evolve over time.
@@
Given a policy $\pi$ and transition function $p$, the occupancy measure $q_{h}^{\pi}$ is defined as the probability of the agent visiting state $s$ and taking action $a$ at step $h$, that is, $q_{h}^{\pi}(s,a;p)=\mathbb{P}\left\{s_{h}=s,a_{h}=a|p,\pi\right\}$ Puterman1994 ; altman1999constrained . The entire collection of these probabilities across all steps, states and actions is denoted by $q^{\pi}(p):=[q_{h}^{\pi}(s,a;p)]_{H\times\mathcal{S}\times\mathcal{A}}$.
@@
We introduce a probability space $(\Omega,\mathcal{F},\mathbb{P})$ to capture the randomness of thresholds, where each $\omega\in\Omega$ is referred to as a threshold ‘event’ (or, scenario), representing a specific realisation of these values. For every constraint $i\in[m]$, the threshold $\alpha_{i}(\omega_{i})$ is drawn from an independent probability space $(\Omega_{i},\mathcal{F}_{i},\mathbb{P}_{i})$. Consequently, the overall threshold vector is $\alpha=(\alpha_{1}(\omega_{1}),\ldots,\alpha_{m}(\omega_{m}))$, where threshold scenario $\omega=(\omega_{1},\ldots,\omega_{m})$. Under the occupancy measure formulation, the optimisation problem then becomes
$\max_{q\in\Delta{(\mathcal{M})}}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha(\omega)),$
where $\omega$ is the threshold scenario. The corresponding random feasible set is
@@ 3.2 Estimators of Thresholds
We then design a Growing-Window (GW) approach to estimating the unknown and stochastic thresholds. In Appendix C, we also give a Monte Carlo (MC) estimator that aggregates all historical data. Compared with the MC estimator, the GW estimator offers better adaptability to environments while preserving asymptotic consistency.
@@
where $N_{h}^{(W_{t}),t}(s,a):=\sum_{l=\max\{1,t-W_{t}+1\}}^{t-1}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ counts visits to state-action pair $(s,a)$ in the last $W_{t}$ episodes before episode $t$. As $t$ increases, the window size $W_{t}$ grows. We prove its asymptotic consistency as follows.
@@
When there is no ambiguity, we may simplify the notation $\hat{\alpha}_{i,h}^{(W_{t}),t}$ to $\hat{\alpha}_{i,h}^{t}$ and $\zeta_{h}^{(W_{t}),t}$ to $\zeta_{h}^{t}$. The proof is given in Appendix C.
@@
With probability $1-\delta$, any possible estimator $\hat{\alpha}_{i}$ satisfies inequality: $\underline{\alpha}_{i}\leq\hat{\alpha}_{i}\leq\overline{\alpha}_{i}$, for any constraint $i\in[m]$. This suggests that pessimistic and optimistic estimators are upper and lower boundaries of all possible estimators, respectively. Therefore, by analysing these pessimistic and optimistic estimators, we can effectively study and bound the performance of the GW estimator.
@@ 3.3 Feasibility and Strong Duality
A policy $\pi$ is defined to be pessimistic if the expected cumulative constraint is no less than the pessimistic threshold $\overline{\alpha}$; i.e., $V_{g}^{\pi}\geq\overline{\alpha}$. Let $\mathcal{F}_{\textup{pes}}$ denote the feasibility set,
@@
Similarly, we define a policy $\pi$ to be optimistic if the expected cumulative constraint is greater than or equal to the optimistic threshold $\underline{\alpha}$, i.e., $V_{g}^{\pi}\geq\underline{\alpha}$, and its feasibility set as below,
@@
From the definitions, pessimistic and optimistic policies are boundaries of all possible policies: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}(\omega)\subseteq\mathcal{F}_{\textup{opt}}$ with probability $1-\delta$.
Pessimistic and optimistic policies can be obtained from the following optimisation problem, respectively.
@@
Each lemma requires the existence of a policy that strictly satisfies the corresponding thresholds. We illustrate these feasible sets under pessimistic, intermediate and optimistic thresholds and their corresponding feasible solutions in Figure 1.
We set the Slater gap under pessimistic and optimistic thresholds as below, respectively,
$\rho:=(r^{T}q^{\pi^{\star}}(p)-r^{T}q^{\pi_{1}})/\min_{i\in[m]}\xi_{i},$
$\rho^{\prime}:=(r^{T}q^{\pi^{\star}}(p)-r^{T}q^{\pi_{2}})/\min_{i\in[m]}\xi^{\prime}_{i},$
@@
As discussed, with probability $1-\delta$, pessimistic thresholds are boundaries for all possible thresholds. If there exists a policy that strictly satisfies pessimistic thresholds, for all possible thresholds, the strong duality holds.
@@ 3.4 A Primal-Dual Approach for CMDPs with Stochastic Thresholds
In this section, we propose SPOT (Stochastic Pessimistic-Optimistic Thresholding), a novel algorithm specifically tailored for CMDPs under stochastic thresholds. Unlike the closest algorithm OptPrimalDual-CMDPefroni2020exploration , which assumes a fixed and known threshold, SPOT tackles the more challenging setting in which the threshold is unknown, and potentially changing over time.
@@
0:  episodes $T$, $\bf\lambda_{1}=0$, stepsizes $\eta_{t}$ and $\eta_{\lambda}$, and initial policy $\pi_{h}^{1}(a|s)=1/A$ for all $s$, $a$, $h$.
@@
9:     Execute $\pi_{t}$ via Algorithm 1 and update counters and empirical model (i.e., $n^{t},\hat{r}^{t},\hat{g}^{t},\hat{p}^{t},\hat{\alpha}^{t}$) via Equation (6)
@@
For a given dual variable, the algorithm first updates the primal policy and subsequently refines the dual value using the estimated value. In the first stage (Line 2), it updates the optimistic estimators $\overline{r}^{t}$ and $\overline{g}^{t}$. Then, it computes the estimator $\alpha_{\diamond}^{t}$ as prescribed in Equation (2). Importantly, depending on the problem context, estimator $\alpha_{\diamond}^{t}$ can be either pessimistic estimator $\overline{\alpha}^{t}$ or optimistic estimator $\underline{\alpha}^{t}$, thereby accommodating different degrees of uncertainty in the thresholds. Thus, at each episode $t$, it yields a new CMDP $\mathcal{M}_{t}:=(\mathcal{S},\mathcal{A},\overline{r}^{t},\overline{g}^{t},\alpha_{\diamond}^{t},\overline{p}^{t})$, which the algorithm then optimises via
@@
where $\overline{G}_{t-1}q^{\pi_{t}}$ is equal to $\hat{V}_{\overline{g}}^{\pi_{t}}$ and the projection set is defined as $\mathcal{C}:=\{\lambda\in\mathbb{R}^{m}:0\leq\lambda\leq\rho\bf 1\}$ with $\rho$ being the Slater gap defined in Section 3.3 and the projection operator $\Pi_{\mathcal{C}}:\mathbb{R}^{m}\to\mathcal{C}$ is defined by $\Pi_{\mathcal{C}}(x):=\arg\min_{y\in\mathcal{C}}\|y-x\|_{2}$. The feasibility proof for Equation (3) is presented in Appendix F. Finally, the policy $\pi_{t}$ is executed to collect new data, which is then used to update the empirical models and threshold estimates for subsequent iterations (Line 9).
@@
  • Stochastic and unknown thresholds vs. fixed thresholds. SPOT can handle thresholds that are unknown and random, going beyond the scope of fixed-threshold methods, which fit in more realistic, critical scenarios.
  • Pessimistic or optimistic thresholds balancing exploration and safety. SPOT allows using either pessimistic or optimistic thresholds, selected according to the application’s risk tolerance: pessimistic thresholds ensure robust feasibility under worst-case assumptions, while optimistic thresholds enable more ambitious exploration when additional risk is acceptable. This twofold approach empowers SPOT to adapt to varying levels of constraint uncertainty.
  • Decent Performance Guarantees. By dynamically refining its threshold estimates, SPOT can maintain performance guarantees that are the same order as fixed-threshold approaches.

### v2

@@ Abstract
This paper studies constrained Markov decision processes (CMDPs) with constraints against stochastic thresholds, aiming at the safety of reinforcement learning in unknown and uncertain environments. We leverage a Growing-Window estimator sampling from interactions with the uncertain and dynamic environment to estimate the thresholds, based on which we design Stochastic Pessimistic-Optimistic Thresholding (SPOT), a novel model-based primal-dual algorithm for multiple constraints against stochastic thresholds. SPOT enables reinforcement learning under both pessimistic and optimistic threshold settings. We prove that our algorithm achieves sublinear regret and constraint violation; i.e., a reward regret of $\tilde{\mathcal{O}}(\sqrt{T})$ while allowing an $\tilde{\mathcal{O}}(\sqrt{T})$ constraint violation over $T$ episodes. The theoretical guarantees show that our algorithm achieves performance comparable to that of an approach relying on fixed and clear thresholds. To the best of our knowledge, SPOT is the first reinforcement learning algorithm that realises theoretical guaranteed performance in an uncertain environment where even thresholds are unknown.
@@ 1 Introduction
Reinforcement learning has enabled major advances in AI, such as AlphaGo silver2016mastering , autonomous vehicle control kendall2019learning , and LLM alignment ouyang2022training , in which an agent is trained to maximise its expected cumulative reward from the environment Sutton1998 . Markov decision processes (MDPs) are widely adopted in reinforcement learning, where transitions and rewards depend only on the current state rather than on the entire history bellman1957markovian . In many applications, ensuring safety is of profound importance – for instance, autonomous vehicles are expected to avoid collisions and adhere to traffic regulations wen2020safe . This motivates constrained Markov decision processes (CMDPs), introducing safety constraints against thresholds into optimisation of the cumulative reward function altman1999constrained .
@@
  • The first approach to safe reinforcement learning agents in such uncertain environment. SPOT leverages a Growing-Window estimator sampling during the agent’s interactions with the environment to estimate the thresholds. Based on this estimator, SPOT employs the primal-dual scheme to integrate the estimated thresholds into the learning process of the agent. SPOT is compatible with two variants that (1) employs pessimistic thresholds to strictly enforce safety, and (2) applies optimistic thresholds to encourage exploration and boost performance. By systematically navigating the latent ‘margin’ between safety and performance without relying on any prior threshold information, our approach offers a dynamic balance between exploitation and exploration.
@@ 2 Preliminaries
An MDP characterises the sequential decision-making process of an agent, positioned in a state $s$ from a finite state space $\mathcal{S}$ (with cardinality $|\mathcal{S}|=S$), which could select an action $a$ from a finite action space $\mathcal{A}$ (with cardinality $|\mathcal{A}|=A$) and then transition to another state $s^{\prime}\in\mathcal{S}$ according to a transition function $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$. The course of each episode starts at an initial state $s_{1}\in\mathcal{S}$ and continues for $H\in\mathbb{N}$ steps, which is named as horizon. At each step $h\in[H]$, the agent receives a non-stationary reward $\tilde{r}_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ drawn from a distribution $\mathcal{R}$ with expectation reward $r_{h}(s,a)$, which means its underlying distribution may vary over time. The agent’s trajectory is determined by a policy $\pi$ consisting of decision rules $(\pi_{1},\ldots,\pi_{H})\in\Pi$, where each $\pi_{h}(\cdot|s)\in\Delta(\mathcal{A})$ is a probability distribution over the action space $\mathcal{A}$. Here, $\Delta(\mathcal{A})$ denotes the probability simplex over the action space $\mathcal{A}$. A Constrained MDP (CMDP) extends an MDP by incorporating $m$ constraints, each defined by a function $g_{i,h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ associated with a threshold $\alpha_{i}\in[0,H]$. These thresholds collectively form a vector $\alpha\in[0,H]^{m}$. In the non-stationary scenario, the immediate constraint $\tilde{g}_{i,h}(s,a)\in[0,1]$ is drawn from a distribution $\mathcal{G}_{i}$ with expectation $\mathbb{E}_{\mathcal{G}_{i}}[\tilde{g}_{i,h}(s,a)]=g_{i,h}(s,a)$ for every $i$-th constraint. Overall, a CMDP can be represented by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},H,p,r,g,\alpha)$.
@@
capture the expected sum of $v$ from step $h$ onward. For brevity, let $V_{v}^{\pi}$ denote $V_{v,1}^{\pi}(s_{1})$, the expected return starting from the initial state $s_{1}$. The primary objective is to find a policy that maximises the expected cumulative reward while satisfying all the constraints; formally,
@@
The reward regret measures the cumulative gap between the rewards obtained by the selected policies and the optimal policy $\pi^{\star}\in\Pi$, and the constraint violation measures how much the constraints are violated during the learning process. The objective of the agent is to select policies $\{\pi_{t}\}_{t=1}^{T}$ such that both $\mathcal{R}_{T}(r)$ and $\mathcal{R}_{T}(g)$ grow sublinearly in the number of episodes $T$.
@@ 3.1 Unknown and Stochastic Environment in Terms of Thresholds
Motivated by the examples above, we propose training a CMDP in which thresholds are unknown, and must therefore be inferred through real-time sampling. Specially, for each constraint $i\in[m]$, the threshold $\tilde{\alpha}_{i,h}(s,a)\in[0,H]$ for any state-action pair $(s,a)$ is drawn from a distribution $\mathcal{L}_{i}$ with support of $[0,H]^{\mathcal{S}\times\mathcal{A}}$, whose expectation is $\mathbb{E}_{\mathcal{L}_{i}}[\tilde{\alpha}_{i,h}(s,a)]=\alpha_{i,h}(s,a)$.
Algorithm 1 depicts how an agent interacts with the environment, when rewards, constraints and thresholds are all stochastic. At each episode $t\in[T]$, the agent chooses a policy $\pi_{t}$, while independently sampling reward $r^{t}\sim\mathcal{R}$, constraint $g_{i}^{t}\sim\mathcal{G}_{i}$ and threshold $\alpha_{i}^{t}\sim\mathcal{L}_{i}$ for the $i$-th constraint. It is important to note that the agent has no prior knowledge of the reward, constraint, and threshold distributions, nor the transition function $p$.
@@
Also motivated by the examples, we note that the environment is not static – rewards, constraints, and thresholds change over time. To take these into account, we define that rewards, constraints and thresholds are drawn from dynamical distributions $\mathcal{R}_{t}$, $\{\mathcal{G}_{i,t}\}_{i=1}^{m}$, $\{\mathcal{L}_{i,t}\}_{i=1}^{m}$ over $T$ episodes, instead of fixed distributions. Consequently, their expectations $r_{h}^{t}(s,a)$, $g_{i,h}^{t}(s,a)$, $\alpha_{i,h}^{t}(s,a)$ also evolve over time.
@@
Given a policy $\pi$ and transition function $p$, the occupancy measure $q_{h}^{\pi}$ is defined as the probability of the agent visiting state $s$ and taking action $a$ at step $h$, that is, $q_{h}^{\pi}(s,a;p)=\mathbb{P}\left\{s_{h}=s,a_{h}=a|p,\pi\right\}$ altman1999constrained ; Puterman1994 . The entire collection of these probabilities across all steps, states and actions is denoted by $q^{\pi}(p):=\{q_{h}^{\pi}(s,a;p)\}_{H\times\mathcal{S}\times\mathcal{A}}$.
@@
We introduce a probability space $(\Omega,\mathcal{F},\mathbb{P})$ to capture the randomness of thresholds, where each outcome $\omega\in\Omega$ represents a specific realisation of these values, referred to as a threshold scenario. For every constraint $i\in[m]$, the threshold $\alpha_{i}(\omega_{i})$ is drawn from an independent probability space $(\Omega_{i},\mathcal{F}_{i},\mathbb{P}_{i})$. Consequently, the overall threshold vector is $\alpha=(\alpha_{1}(\omega_{1}),\ldots,\alpha_{m}(\omega_{m}))$, where the threshold scenario $\omega=(\omega_{1},\ldots,\omega_{m})$. Under the occupancy measure formulation, the optimisation problem then becomes
$\max_{q\in\Delta{(\mathcal{M})}}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha(\omega)).$
where $G$ denotes the constraint matrix induced by the values $g_{i,h}(s,a)$ for all $i\in[m]$, $h\in[H]$, and $(s,a)\in\mathcal{S}\times\mathcal{A}$. Moreover, the corresponding stochastic feasible set is
@@ 3.2 Estimators of Thresholds
We then design a Growing-Window (GW) approach to estimating the unknown and stochastic thresholds. Unlike the Monte Carlo (MC) estimator that aggregates all historical data (see Appendix C), the GW estimator dynamically focuses on more recent observations, offering better adaptability to non-stationary environments while still preserving asymptotic consistency.
@@
where $N_{h}^{(W_{t}),t}(s,a):=\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ counts visits to state-action pair $(s,a)$ in the last $W_{t}$ episodes before episode $t$. As $t$ increases, the window size $W_{t}$ grows. We prove its asymptotic consistency as follows.
@@
When there is no ambiguity, we simplify the notation $\hat{\alpha}_{i,h}^{(W_{t}),t}$ to $\hat{\alpha}_{i,h}^{t}$ and $\zeta_{h}^{(W_{t}),t}$ to $\zeta_{h}^{t}$. The proof is given in Appendix C.
@@
With probability at least $1-\delta$, any possible estimator $\hat{\alpha}_{i}$ satisfies inequality: $\underline{\alpha}_{i}\leq\hat{\alpha}_{i}\leq\overline{\alpha}_{i}$, for any constraint $i\in[m]$. This suggests that pessimistic and optimistic estimators are upper and lower boundaries of all possible estimators, respectively. Therefore, by analysing these pessimistic and optimistic estimators, we can effectively study and bound the performance of the GW estimator.
@@ 3.3 Feasibility and Strong Duality
A policy $\pi$ is defined to be pessimistic if the expected cumulative constraint is no less than the pessimistic threshold; i.e., $V_{g}^{\pi}\geq\overline{\alpha}$. Let $\mathcal{F}_{\textup{pes}}$ denote its feasibility set,
@@
Similarly, we define a policy $\pi$ to be optimistic if the expected cumulative constraint is greater than or equal to the optimistic threshold, i.e., $V_{g}^{\pi}\geq\underline{\alpha}$, and its corresponding feasibility set as below,
@@
From the definitions, pessimistic and optimistic policies are boundaries of all possible policies: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}(\omega)\subseteq\mathcal{F}_{\textup{opt}}$ with probability at least $1-\delta$.
Pessimistic and optimistic policies can be obtained from the following optimisation problems, respectively,
@@
Each lemma requires the existence of a policy that strictly satisfies the corresponding thresholds. We illustrate these feasible sets and their corresponding feasible solutions under pessimistic, intermediate (thresholds derived from random sampling), and optimistic thresholds in Figure 1.
We set the Slater gaps under pessimistic and optimistic thresholds as follows, respectively,
$\rho:=(r^{T}q^{\pi^{\star}}-r^{T}q^{\pi_{1}})/\min_{i\in[m]}\xi_{i},$
$\rho^{\prime}:=(r^{T}q^{\pi^{\star}}-r^{T}q^{\pi_{2}})/\min_{i\in[m]}\xi^{\prime}_{i}.$
@@
As discussed, with probability at least $1-\delta$, pessimistic thresholds are upper bounds for all possible thresholds. If there exists a policy that strictly satisfies pessimistic thresholds, then it also satisfies all possible thresholds, ensuring that the strong duality holds.
@@ 3.4 A Primal-Dual Approach for CMDPs with Stochastic Thresholds
In this section, we propose SPOT (Stochastic Pessimistic-Optimistic Thresholding), a novel algorithm specifically tailored for CMDPs with stochastic thresholds. Unlike the closest algorithm OptPrimalDual-CMDPefroni2020exploration , which relies on fixed and fully known thresholds, SPOT is designed to address the more challenging and realistic scenarios where thresholds are both unknown and dynamically evolving over time.
@@
0:  number of episodes $T$, $\bf\lambda_{1}=0$, stepsizes $\eta_{t}$ and $\eta_{\lambda}$, and initial policy $\pi_{h}^{1}(a|s)=1/A$ for all $s$, $a$, $h$.
@@
9:      Execute $\pi_{t}$ via Algorithm 1 and update counters and empirical model (i.e., $\hat{\alpha}^{t},\hat{r}^{t},\hat{g}^{t},\hat{p}^{t}$) via Equations (2) and (7)
@@
For a given dual variable, the algorithm first updates the primal policy based on the estimated $Q$-values and subsequently refines the dual value using the estimated value. In the first stage (Line 2), it updates the optimistic estimators $\overline{r}^{t}$, $\overline{g}^{t}$ and $\overline{p}^{t}$. Then, it computes the estimator $\alpha_{\diamond}^{t}$ as prescribed in Equation (3). Importantly, depending on the problem context, estimator $\alpha_{\diamond}^{t}$ can be either pessimistic estimator $\overline{\alpha}^{t}$ or optimistic estimator $\underline{\alpha}^{t}$, thereby accommodating different degrees of uncertainty in the thresholds. Thus, at each episode $t$, it yields a new CMDP $\mathcal{M}_{t}:=(\mathcal{S},\mathcal{A},\overline{r}^{t},\overline{g}^{t},\alpha_{\diamond}^{t},\overline{p}^{t})$, which the algorithm then optimises via
@@
where $\overline{G}_{t-1}q^{\pi_{t}}$ corresponds to the value estimate $\hat{V}_{\overline{g}}^{\pi_{t}}$ and the projection set is defined as $\mathcal{C}:=\{\lambda\in\mathbb{R}^{m}:0\leq\lambda\leq\rho\bf 1\}$ with $\rho$ being the Slater gap defined in Section 3.3 and the projection operator $\Pi_{\mathcal{C}}:\mathbb{R}^{m}\to\mathcal{C}$ is defined by $\Pi_{\mathcal{C}}(x):=\arg\min_{y\in\mathcal{C}}\|y-x\|_{2}$. The feasibility proof for Equation (4) is presented in Appendix F. Finally, the policy $\pi_{t}$ is executed to collect new data, which is then used to update the threshold estimates and empirical models for subsequent iterations (Line 9).
@@
  • Stochastic and unknown thresholds vs. fixed thresholds. SPOT can handle thresholds that are unknown and random, going beyond the scope of fixed-threshold methods, which fit in more realistic scenarios.
  • Dual-mode thresholding for exploration and safety. SPOT introduces a novel dual-mode thresholding mechanism, enabling the selection of either pessimistic or optimistic thresholds based on the application’s risk tolerance: pessimistic thresholds ensure robust feasibility under worst-case assumptions, while optimistic thresholds enable more aggressive exploration when additional risk is acceptable. This twofold approach empowers SPOT to adapt to varying levels of constraint uncertainty.
  • Strong Performance Guarantees. By continuously refining its threshold estimates, SPOT can maintain performance guarantees that are the same order as fixed-threshold approaches, while maintaining adaptability to dynamic changes.